Consider the following Decision alternative for the Raman Pahwa, he wants to invest in

stocks, and thought about two situations about tomorrow’s market condition. The figures

(in INR) in the following table exhibit profit per unit of stock-investment.

payoff table:

  • Favorable market

Lakshmi pvt ltd 55

Mehta Groups of industries 43

Surya 29

LT energy 15

  • Unfavorable market

Lakshmi pvt ltd 26

Mehta Groups of industries 38

Surya 43

LT energy 51


  1. Draw the decision tree

2. If we assign the following probabilities to the states of nature, then determine the

EMV decision.

  1. P(s1) = .4 P(s2) = .1 P(s3) = .3 P(s4) = .2
1
Expert's answer
2022-03-01T06:03:02-0500



EMV decision = (4*2) + (4*3) = 8+12= 20